Circular No. 8371

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1186 SCORPII = NOVA SCORPII 2004 N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V1186 Sco has been given to this nova (cf. IAUC 8369).

E. Waagen, AAVSO, reports that V1186 Sco has brightened, as indicated by the following magnitude estimates (visual unless otherwise noted): July 6.413 UT, 10.8 (J. Bedient, Honolulu, HI); 7.358, 10.6 (M. Linnolt, Honolulu, HI); 8.064, 9.9 (B. Monard, Tiegerpoort, South Africa; unfiltered CCD); 8.175, V = 10.2 (R. James, Las Cruces, NM; CCD); 8.440, 10.3 (Linnolt); 8.79, 9.5 (Monard; unfiltered CCD); 8.82, B = 11.0 (Monard; CCD); 9.331, 9.8 (Linnolt); 9.489, 9.7 (R. Axelson, Chapel Hill, Qld., Australia).

SUPERNOVA 2004da IN NGC 6901 C. L. Gerardy and G. H. Marion, University of Texas at Austin, report

that a low-resolution optical spectrogram (resolution 300; range 410–1000 nm) of SN 2004da (cf. IAUC 8370), obtained on July 8.26 UT with the 9.2-m Hobby/Eberly Telescope (+ Marcario Low-Resolution Spectrograph) by M. Shetrone and F. Deglman, indicates that it is a peculiar type-Ia supernova, a few days after maximum light. The spectrum resembles that of SN 1997br when eight days past maximum light (Li et al. 1999, A.J. 117, 2709), albeit exhibiting somewhat larger expansion velocities, perhaps indicating that SN 2004da is at an earlier epoch. Adopting the redshift of NGC 6901 from the NASA/IPAC Extragalactic Database (4794 km/s), the expansion velocity from the absorption minimum of the Si II feature (rest wavelength 635.5 nm) is 15600 km/s.

SUPERNOVA 2003lz IN ESO 428-G13 Y. Chen, Z. Lin, K. Huang, H. Lin, and W. Ip, Institute of Astronomy,

National Central University (note that “Taiwan” was spelled incorrectly on IAUC 8364); and Y. Qiu and J. Hu, National Astronomical Observatories of China, report the discovery of an apparent supernova (mag 17.1) on unfiltered archival CCD images taken on 2003 Oct. 30.83 UT with the 1-m Lulin Observatory telescope. The new object was located at α = 7^{h}16^{m}25^{s}.24, δ = −29^{o}37′04′′.5 (equinox 2000.0), which is 14′′.3 west and 10′′.3 north of the center of ESO 428-G13. Additional unfiltered (red) magnitudes for SN 2003lz: 2003 Nov. 6.83, 16.5; Dec. 7.77, 17.9; 9.83, 18.0; 19.83, 18.8; 2004 Jan. 5.83, [19.0; Apr. 26.50, [19.0. Nothing is visible at the location of SN 2003lz on Palomar Sky Survey images.

2004 July 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green